# Semi-decentralized Training of Spatio-Temporal Graph Neural Networks for Traffic Prediction

### v1

@@ VIII Conclusion
Experimental results on the METR-LA and PeMS-BAY dataset show that semi-decentralized setups are comparable to centralized approaches in performance metrics, while offering advantages in terms of scalability and fault tolerance. We also highlight often overlooked issues in existing literature for distributed ST-GNNs, such as the variation in model performance across different geographical areas due to region-specific traffic patterns, and the significant communication overhead and computational costs that arise from the large receptive field of GNNs, leading to substantial data transfers and increased computation of partial embeddings.

### v2

@@ VIII Conclusion
Experimental results on the METR-LA and PeMS-BAY datasets show that semi-decentralized setups are comparable to centralized approaches in performance metrics, while offering advantages in terms of scalability and fault tolerance. We also highlight often overlooked issues in existing literature for distributed ST-GNNs, such as the variation in model performance across different geographical areas due to region-specific traffic patterns, and the significant communication overhead and computational costs that arise from the large receptive field of GNNs, leading to substantial data transfers and increased computation of partial embeddings. However, due to the planar nature of graphs, cloudlets costs remain fixed as the network grows, unlike the growing costs in a centralized approach.